Solve the differential equation 𝑑𝑦 𝑑𝑥 + 2𝑦 tan 𝑥 = sin 𝑥 , 𝑦 ( 𝜋 3 ) = 0.


1
Expert's answer
2021-10-11T16:28:59-0400

Let us solve the differential equation "\frac{dy}{dx}+2y\tan x=\sin x,\ y(\frac{\pi}3)=0." Let us divide both parts by "\cos^2 x." Then we get the differential equation "\frac{1}{\cos^2 x}y'+\frac{2\sin x}{\cos^3 x}y=\frac{\sin x}{\cos^2 x}," which is equivalent to "(\frac{1}{\cos^2 x}y)'=\frac{\sin x}{\cos^2 x}." It follows that

"\frac{1}{\cos^2 x}y=\int\frac{\sin x}{\cos^2 x}dx=-\int\frac{d(\cos x)}{\cos^2 x}=\frac{1}{\cos x}+C," and hence

"y=\cos x+C\cos^2 x" is the general solution of the differential equation.

Since "y(\frac{\pi}3)=0," we get that "0=\cos\frac{\pi}3+C\cos^2\frac{\pi}3=\frac{1}2+C\frac{1}4." Consequently, "C=-2." We conclude that the solution of "\frac{dy}{dx}+2y\tan x=\sin x,\ y(\frac{\pi}3)=0," is

"y=\cos x-2\cos^2 x."